OpenStudy (anonymous):

solve the system by graphing -2x=2y-4 2x-y=-5

OpenStudy (anonymous):

@piglet9

OpenStudy (anonymous):

Put equation in the slope intercept form (y = mx +b) by solving the equation for y

OpenStudy (jdoe0001):

ditto

OpenStudy (anonymous):

2x+2y=4 or x+y=2 when x=0,y=2 when y=0,x=2 draw line AB y=2x+5 when x=0,y=5 when x=2,y=9 draw line CD Draw the graphs and see where the two lines AB and CD intersect.

OpenStudy (anonymous):

this is not any of the answers listed.

OpenStudy (anonymous):

this is not the answer. you draw on the graph lines AB and CD ,then the point of intersection is the answer.

OpenStudy (anonymous):